OpenStudy (anonymous):

Decide if the following problem is an example of a permutation or a combination. State your rationale. For a study, 4 people are chosen at random from a group of 10 people. How many ways can this be done? a. combination - set of objects arranged in a certain order c. combination - an unordered arrangement of objects b. permutation - set of objects arranged in a certain order d. permutation - an unordered arrangement of objects

OpenStudy (amistre64):

P, permutation, picking, you pick an order C, looks like a G, combinations, you group things so order is orrelevant

OpenStudy (amistre64):

**irrelevant

OpenStudy (anonymous):

okay, thank you. Could you help me out with some more permutation questions?

OpenStudy (amistre64):

maybe, just keep that rule of thumb in mind

OpenStudy (anonymous):

How would you solve this?

OpenStudy (amistre64):

how many groups of 1, can you make out of 325 elements?

OpenStudy (anonymous):

Oh that's how you read that? Okay, that makes sense. So it'd be 325

OpenStudy (amistre64):

yep, but let me dbl chk myself spose we had 3 elements a b c how many groups of 1 can we make? a b c yep 3 groups of 1

OpenStudy (anonymous):

Okay! how would you solve a more complicated one like this?

OpenStudy (amistre64):

do you know what a factorial is?

OpenStudy (anonymous):

Okay! how would you solve a more complicated one like this?

OpenStudy (amistre64):

or, we can use the identity: k! nCk = nPk nCk = nPk/k! k! is notation for a factorial number

OpenStudy (anonymous):

It's where you mutiply descending numbers?

OpenStudy (anonymous):

*multiply. For ex) 3!= 3 x 2 x 1

OpenStudy (amistre64):

yes, and nPk means we descend for k terms \[10C4*4C2\] \[\frac{10P4}{4!}*\frac{4P2}{2!}\] \[\frac{10.9.8.7}{4.3.2}*\frac{4.3}{2}\] \[\frac{10.9.8.7}{8.3}*\frac{4.3}{2}\] \[\frac{10.9.7}{3}*\frac{4.3}{2}\] \[10.3.7*2.3\]

OpenStudy (amistre64):

nPn = n!

OpenStudy (anonymous):

oh okay, I think I get it now

OpenStudy (amistre64):

since nCk = (nPk)/k! and k! = kPk \[nCn=\frac{nPn}{nPn}\] and anything divided by itself is 1, except for 0/0 so, what this simply means is that there is only 1 way to group a set n, by size n